OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2

OpenStudy (anonymous):

ok i think you use the formula y=mx+b

OpenStudy (anonymous):

yes you do

OpenStudy (anonymous):

How do you set up the equation?

OpenStudy (anonymous):

1=m4+b 2=m5+b

OpenStudy (anonymous):

y=1x-3